BRICS Research Series, 2007

June 27, 2008

This document is also available as PostScript and DVI.

References

[RS-07-18] Abstract, PostScript, PDF, DVI.

Jan Midtgaard. Control-Flow Analysis of Functional Programs. December 2007. iii+38 pp.

[RS-07-17] Luca Aceto, Willem Jan Fokkink, and Anna Ingólfsdóttir. A Cancellation Theorem for 7BCCSP. December 2007. 30 pp.

[RS-07-16] Abstract, PostScript, PDF, DVI.

Olivier Danvy and Kevin Millikin. On the Equivalence between Small-Step and Big-Step Abstract Machines: A Simple Application of Lightweight Fusion. November 2007. ii+11 pp. To appear in Information Processing Letters (extended version). Supersedes BRICS RS-07-8.

[RS-07-15] Jooyong Lee. A Case for Dynamic Reverse-code Generation. August 2007. ii+10 pp.

[RS-07-14] Abstract, PostScript, PDF, DVI.

Olivier Danvy and Michael Spivey. On Barron and Strachey’s Cartesian Product Function. July 2007. ii+14 pp.

[RS-07-13] Martin Lange. Temporal Logics Beyond Regularity. July 2007. 82 pp.

[RS-07-12] Gerth Stølting Brodal, Rolf Fagerberg, Allan Grønlund Jørgensen, Gabriel Moruz, and Thomas Mølhave. Optimal Resilient Dynamic Dictionaries. July 2007.

[RS-07-11] Abstract, PostScript, PDF, DVI.

Luca Aceto and Anna Ingólfsdóttir. The Saga of the Axiomatization of Parallel Composition. June 2007. 15 pp. To appear in the Proceedings of CONCUR 2007, the 18th International Conference on Concurrency Theory (Lisbon, Portugal, September 4–7, 2007), Lecture Notes in Computer Science, Springer-Verlag, 2007.

[RS-07-10] Abstract, PostScript, PDF.

Claus Brabrand, Robert Giegerich, and Anders Møller. Analyzing Ambiguity of Context-Free Grammars. May 2007. 17 pp. Full version of paper presented at CIAA ’07.

[RS-07-9] Janus Dam Nielsen and Michael I. Schwartzbach. The SMCL Language Specification. March 2007.

[RS-07-8] Abstract, PostScript, PDF, DVI.

Olivier Danvy and Kevin Millikin. A Simple Application of Lightweight Fusion to Proving the Equivalence of Abstract Machines. March 2007. ii+6 pp.

[RS-07-7] Abstract, PostScript, PDF, DVI.

Olivier Danvy and Kevin Millikin. Refunctionalization at Work. March 2007. ii+16 pp. Invited talk at the 8th International Conference on Mathematics of Program Construction, MPC ’06.

[RS-07-6] Abstract, PostScript, PDF, DVI.

Olivier Danvy, Kevin Millikin, and Lasse R. Nielsen. On One-Pass CPS Transformations. March 2007. ii+19 pp. Theoretical Pearl to appear in the Journal of Functional Programming. Revised version of BRICS RS-02-3.

[RS-07-5] Abstract, PostScript, PDF, DVI.

Luca Aceto, Silvio Capobianco, and Anna Ingólfsdóttir. On the Existence of a Finite Base for Complete Trace Equivalence over BPA with Interrupt. February 2007. 26 pp.

[RS-07-4] Abstract, PostScript, PDF, DVI.

Kristian Støvring and Søren B. Lassen. A Complete, Co-Inductive Syntactic Theory of Sequential Control and State. February 2007. 36 pp. Appears in the proceedings of POPL 2007, p. 161–172.

[RS-07-3] Abstract, PostScript, PDF.

Luca Aceto, Willem Jan Fokkink, and Anna Ingólfsdóttir. Ready To Preorder: Get Your BCCSP Axiomatization for Free! February 2007. 37 pp.

[RS-07-2] Abstract, PostScript, PDF, DVI.

Luca Aceto and Anna Ingólfsdóttir. Characteristic Formulae: From Automata to Logic. January 2007. 18 pp.

[RS-07-1] Abstract, PostScript, PDF.

Daniel Andersson. HIROIMONO is NP-complete. January 2007. 8 pp.